OpenStudy (anonymous):

A pineapple was found to have two sets of opposing spirals formed by its leaves. The clockwise spiral has x leaves. The number of leaves in the counterclockwise direction is 5 more than that in the clockwise direction. These leaf numbers are consecutive numbers in the Fibonacci sequence. which equation is true for the number of leaves in the two spirals? x+5/5=1.6 x/5=1.6 5/x=1.6 x+5/x=1.6

Directrix (directrix):

Would you click on the "Draw" tab and write the following: x+5/5=1.6 x/5=1.6 5/x=1.6 x+5/x=1.6 As is, I cannot determine which terms are in the numerator and which in the denominator. Thanks.